OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 - x - 20

OpenStudy (darkprince14):

Since the last value was a negative number (-20), it means that the factors have different operations.. (x-b)(x+a) = x2 - x - 20 you need to find b > a and b-a = 1 and b*a = 20..